OpenStudy (favorite):

hi! this looks like a VERY basic algebra question but it something different! pwease help ಠ ∩ ಠ Teachers told us to learn the value of x but i'm curious and wanna know how to find it.. (this calculation is actually related to finding \(p^H\) of dilute acids )

OpenStudy (favorite):

\(10^{-14}=(x+10^{-7})x\) find \(x\) without using any calculator

OpenStudy (favorite):

i think that quadratic formula can't help us with this one..

OpenStudy (mathstudent55):

This is a quadratic equation in x. You can use the quadratic formula. Start by multiplying out the right side using the distributive property.

OpenStudy (favorite):

but that becomes very lengthy tho

OpenStudy (mathstudent55):

Not at all. It's just a quadratic with 3 terms.

OpenStudy (tkhunny):

Why does that matter? If it's a solution, then go find it. (It is NOT lengthy.)

OpenStudy (mathstudent55):

Copy the left side. Then distribute x on the right side. Can you do it and show what you got?

OpenStudy (tkhunny):

First off, just look at it. x = 0 is ALMOST a solution.

OpenStudy (favorite):

yeah i'm typing

OpenStudy (favorite):

\(10^{-14}=(x+10^{-7})x\) \(x^2+10^{-7}x-10^{-14}=0\) \(x=\Large \frac{-10^{-7} \pm\sqrt{1\color{red}{-4\times (-10^{-14})}}}{2}\)

OpenStudy (favorite):

should i ignore the red part? it is almost 0 (we need approximate solutions but not 0)

OpenStudy (tkhunny):

Not quite. Check that "1" under the radical.

OpenStudy (mathstudent55):

In the root, there should be b^2. You have 1.

OpenStudy (favorite):

oops *\(x=\Large \frac{-10^{-7} \pm\sqrt{10^{-14}\color{red}{-4\times (-10^{-14})}}}{2}\)

OpenStudy (mathstudent55):

\(\Large x= \frac{-10^{-7} \pm\sqrt{(10^{-7})^2 -4\times (-10^{-14})}}{2}\)

OpenStudy (mathstudent55):

Yes, good. Now simplify the radical.

OpenStudy (favorite):

\(x=\Large \frac{-10^{-7} \pm\sqrt{10^{-14}\color{red}{+4\times (10^{-14})}}}{2}\) \(x=\Large \frac{-10^{-7} \pm\sqrt{5\times (1{0^{-14})}}}{2}\) \(x=\Large \frac{-10^{-7} \pm10^{-7}\sqrt{5}}{2}\)

OpenStudy (mathstudent55):

\(\Large x= \frac{-10^{-7} \pm\sqrt{5 \times 10^{-14}}}{2}\) \(\Large x= \frac{-10^{-7} \pm\sqrt{5} \times 10^{-7}}{2}\) Good.

OpenStudy (favorite):

\(x=\Large \frac{-10^{-7} + 10^{-7}\sqrt{5}}{2}\) (x is positive ->given) \(x=\Large \frac{10^{-7}( \sqrt{5}-1) }{2}\) \(x=0.618 \times 10^{-7}\)

OpenStudy (favorite):

thanks! :) this is the value of x I was looking for! i think its better to learn the value because it is used very often in these pH calculations

OpenStudy (mathstudent55):

\(x = 6.18 \times 10^{-8} \)

OpenStudy (mathstudent55):

Great. You're welcome.